LUCIFER Commissioning - 2008 December 12 UT

Observers: W. Seifert, N. Ageorges, M. Jutte, J. Hill (on mtn)
Telescope Operator: W. Wack
Software Support: P. Grenz (in Tucson), V. Knierim (in Germany)

Summary

Night starts with lightly scattered cirrus and slowly builds up to heavy cirrus. Seeing was good the whole time seen as 0.32 to 0.65 arcsec FWHM on the guider. Thermal management of the oil is better with some new manual tricks we learned. This may have contributed to improved seeing. Occasional problems with GCS and PCS getting confused. We lost about an hour to assorted small problems.

Still have issues with primary jumping up to 1 arcsec when it gets an active correction, (but not when its collimation is updated). ??

Did alot of tests with offsetting from LUCIFER. Took some data on NGC 1569 but it suffered from clouds and some image quality issues.

Details (times in UT below)

Preparation

New patches of OSS and GCS.

The day guys replaced failed cooling pump #2. J. Kern and J. Hill have discovered a new trick for clearing ice from the HBS heat exchanger (by pulsing the glycol ball valve which seems to flex the heat exchanger and loosen the ice). We have managed to regulate the tank temperature to something approaching outside ambient without freezing up the heat exchanger.

Open

23:22 Open Dome - T=4 degC, sky partly cloudy with scattered cirrus, D = -6.6 degC, LBT wind 10 m/sec looking very laminar.

Had to restart oacserver again to get communications with AGW working. We restarted by rebooting agw-control.

00:45 Adjust CV0423 to 50% open, and the manual ball valve upstairs is also 40% open.

00:54 Rollback again to the previous version of GCS 1.9.51, as today's new version had trouble talking to AGW.

01:00 Preset in parallactic mode. IE = -6.6 CA = -0.2

01:14 Preset failed because star was too far out of focus. Wayne had cleared active optics but not secondary offload. However, that wasn't the problem. Or was it?

01:17 Preset succeeds but horrible coma. We don't understand what happened here. Maybe we didn't press collimate after removing the secondary offload vector?

01:30 Active optics has converged, and the images on the guider are a lovely 0.4 arcsec FWHM with a hint of elongation.

Checking LUCIFER image quality

01:45 Walter is collimating with several different guide stars to look at the image quality.

01:50 Adjust the HBS Ta force on WinCC down from 11.80 (18) to 11.50 (15) to get pressure of 122 Bar.

01:54 Input oil +1.6 Output oil -5.8 Pads are about 6 degC

01:54 Preset failed because couldn't find star. in acquire000167 (really wasn't there).

01:55 Preset TRACK to check pointing. IE = +7.6 CA = -0.2

01:59 Preset TRACK

02:01 Preset ACTIVE

02:10 We saw a focus shift of 230 nm when changing from one guidestar position to the next. Small, but noticeable in this good seeing.

02:19 The image still jumps when we update the primary, but the non-primary jumps seem noticably smaller than last night. The WFS is also significantly better than last night getting down to 150 nm rms.

02:24 Manually sent Z22 -20 nm which is what WFS has been asking for.

02:25 Preset ACTIVE 21:04 +30 AZ=-78 EL=46

Checking LUCIFER focus

Integrated 2x15 sec exposures, probe is at x=5.68 y=88.23 mm

luci_20081211_0021 27.7 (corresponding to 31.6 on-axis)
luci_20081211_0222 21.7
luci_20081211_0223 24.7 (seeing looks worse)
luci_20081211_0224 27.7
luci_20081211_0225 30.7
luci_20081211_0226 33.7
luci_20081211_0227 18.7 (vibration?)
luci_20081211_0228 27.7

Focus conclusion is that the nominal focus (31.6 mm on-axis) looks OK.

Note that the WFS seems to be only getting ~60% (should be 80%) of the Z4 correction when the defocus is large (1500 nm). The GCS scale factor in use is -1.5.

Pause

This log takes a break while John has a shower. I think Walter was measuring the center of rotation on LUCIFER.

Resume

03:56 We went upstairs to check the cooling oil flow. There was a low but steady flow out of the weep hole. We cycled the ball valve a few times without having much effect on the flow. Open manual ball valve 100% and close CV0423 to 40%. Trying to regulate input oil temp near 0 degC without freezing up the heat exchanger.

04:27 Open manual ball valve and close CV0423 to 42%.

Checking XY offsets in the LUCIFER focal plane

04:37 We are just cruising along with guider images between 0.40 and 0.45 arcsec. WFS convergence is down around 80 nm rms. We see jumps when the primary updates. And we see jumps of not more than 0.2 arcsec during normal guiding. I think we must conclude that many of the small 0.4 sec jumps seen last night were the result of local seeing due to the HBS pads being overly warm.

04:39 Input oil +0.6 Output oil -8.9, Pads in 5, Pads out 7

04:40 Adjusted Ta force down to 11.40 (14).

04:56 Set Z11 gain on M1 to zero.

05:00 I was able to confirm my previous result from the IRTC observing. Repeated pressing of the M1 collimate button (say at 20 sec intervals) does not produce any obvious jumps of the image. So then we ask, how is that different from sending a tiny focus correction associated with Z11 at 1 minute intervals while WFSing? One difference is that the collimation updates from the Elevation look-up table are always monotonic.

05:05 Set Z11 gain on M1 back to 1.

05:08 Set active corrections to go the primary. We still get a half arcsec jump when WFS applies X=-0.016 Y=-0.024 RX=-0.258 RY=0.169. The next correction gave a total of X=0.026 Y=-0.026 RX=-0.281 RY=-0.282. Perhaps this indicates that non-monotonic corrections are influencing backlash in the hardpoints.

05:12 Preset fails when it can't find the star. Perhaps because we are at EL=86? Telescope went to PA=0, but pcs.side[0].pointingStatus.positionAngle reads 1.5707 in the Data Dictionary.

05:20 Preset ACTIVE to a lower star with PA=0. The preset succeeded. But DD still reports 90 in DDViewer launched from to-station.

05:24 Stop and Restart PCS.

05:25 Preset ACTIVE with PA=3.54 deg.

05:26 Preset ACTIVE with PA=12 deg.

Dec 11 22:26:40 lbtmu102 LBT_PCS: setNewTarget. 53.0662 0 01:49:46.950000000 +48:37:53.400000000 Mode:POSITION Angle:12 0
but DDViewer still says 0.

Just before 03:47:48 Was a preset with non-zero PA that worked.

05:38 Stopped PCS again, and verified that the exacutable disappeared. Still didn't update DD. See new Issue #1894.

06:00 Send active corrections to "default"

06:02 Nancy reports that the offset commands and absorb are not doing what she expects them to do.

Note that GSC was happily WFSing on a binary star for the last 10 minutes or so.

06:05 Input oil -0.7 Output oil -9.1

06:07 Set Z11 gain on M1 to 0.01. And set Secondary Offload to no. This should update M1 collimation without applying any other position corrections.

06:07 Set Z11 gain on M1 to 0.1 (because Chris was too clever to accept tiny adjustments)

06:20 LUCIFER images are showing defocus and elongation while the guider has round images in the 0.45 to 0.5 arcsec range. Could we have jumps in the Z-axis? LUCIFER focus seems better on right side of chip. Could we have residual binodal astigmatism?

06:26 Back to default active optics mode. Corrections to "default", Secondary Offload = yes, Hold WFS Corrections = no, Z11 gain = 1.

06:33 NGC 1569 RA=04:30 DEC +65, AZ=-1:30 EL=57.8 Seeing on guider is 0.5 arcsec. Polaris monitor reports 1.25

06:45 Guidestar was reported lost. But it was only the the script offsetting out of the patrol field.

06:58 Input oil -0.9 Output oil -9.1

07:05 Slew to zenith to check LUCIFER and insert pinhole mask. Pinhole mask shows LUCIFER images 0.9 x 1.1 pixels. Slight degradation on left edge up to 2.5 pixels due to detector tilt.

07:57 Preset TRACK

08:01 Preset ACTIVE couldn't find star IE=+7.9 CA=+3.0 (We didn't reset them after restarting PCS.)

08:07 Why is guide star significantly out of focus when we have been stopped for an hour? Possible clouds.

08:10 WFS continues, but both GCSGUIs are frozen. Killed and restarted John's GUI, and Wayne's GUI recovers.

08:15 M1 hits -Y limit at -2.6 mm. Not sure I understand why this happened at EL=53. Increase global offset by +1.0 for both M1 and M2. Lost 5 minutes.

08:25 Kill both GCSGUIs again. Restart only Wayne's and it is working.

08:36 Wayne's GCSGUI has frozen again. We restarted PCSGUI and PRESET GUI along the way.

08:59 LUCIFER sends an offet command but it never comes back

09:15 Input oil -1.3 Output oil -9.7

09:20 Kill and restart GCS. The message emitted at kill time was Invalid Buffer RPC error. This also cleared up the GCSGUI.

09:21 Preset TRACK

09:28 rotator stopped tracking - MCSPU complaining of bad polynomials
Dec 12 02:33:26 lbtmu102 LBT_PCS: ROT RPC Error: 000000B000

09:32 stop and restart PCS. Polynomials are good again. Lost 10 minutes.

09:33 preset ACTIVE

09:45 WFS claims to be at 94 nm rms , but guidestar looks a bit comatic (or has a faint companion)

LUCIFER software testing

09:46 preset ACTIVE - Talking to Volker about offsetting commands in LUCIFER.

09:50 Moonlit clouds are getting heavy enough that they are interfering with WFS. Seeing on guider still 0.5 arcsec.

09:59 preset ACTIVE to a bright guidestar to burn through the cirrus

10:07 Clear C00 to get rid of the double image that built up when correcting on clouds. Apparently it was a 2 arcsec binary making fake astigmatism.

10:08 preset ACTIVE to a non-binary guidestar. EL=35 Didn't find any star in acquisition.

10:16 preset ACTIVE to 9th mag guidestar. Sky now has 2-4 mags of cirrus. Probe couldn't move to position - opposite PA.

10:17 preset ACTIVE - succeeded with star at edge of acq field. This star is too bright for WFS, so manually adjust focus until a thicker cloud drifts over.

10:26 edit AIP_L.cfg to allow WFS from 10 to 40 sec, and guide from 0.2 to 4.5 sec. Set threshold for correction of Z22 to 100 nm rms.

10:35 Clouds are thick enough that the 9th mag star is working OK now.

11:10 Input oil -1.3 Output oil -9.8 Pads In 4, Pads Out 5, Ambient=4

11:12 Seeing on guider is down to 0.32 arcsec in 0.5 sec exposure. Too bad it has 3 mags of cirrus in front.

11:58 modify lbt.conf HBThermalLoopOffset float -7.0
to account for the glycol return temp -3.5 that we get when holding Input oil at -1.3 for ambient 4 degC.

12:00 Stop and restart ECS to pickup the new parameters.

12:10 Experimenting with pressing the PSF M1 Collimate button every 20-30 sec, since WFS isn't sending many corrections on this bright star. The frequent update of the Collimation doesn't produce any big (1 arcsec) jumps, although sometimes it makes a small (0.3 arcsec) jump. Maybe we should send a steady stream of stepFocus commands? (That would screw up the guide and WFS loops.)

12:17 Input oil -1.2 Output oil -8.9 CV0423 in auto at 41.3% and TT0423 is at -2.46.

12:18 Preset ACTIVE with mag 10 guidestar.

12:50 Input oil -2.2 Output oil -10.3 CV0423 in auto at 42.8% and TT0423 is at -3.79.

13:05 Guiding has stopped with saturation in 200 msec exposure (a clear patch), so star has drifted out of hotspot. We need WFS to recognize this and not send bogus corrections. (The problem is how to differentiate poor collimation from a miscentered star.)

Dec 12 06:16:03 lbtmu107 LBT_left_GCS: error requesting an image in GuidingThread. Lost connection?

13:18 Kill and restart GCS to see if it fixes the above problem.

13:22 Primary into -Y limit again because of some bogus corrections.

13:23 Preset ACTIVE couldn't find star. Finish clearing all the active corrections.

13:24 Preset ACTIVE converges quickly with some focus help at the beginning.

13:47 Observing ends chasing an obscure error in the LUCIFER software into the sunrise.

Close

13:50 Close Dome T=2.7degC, D=-10 degC. LBTO wind 4 m/sec, thick cirrus all around.

Supplementary Collimation Data

At Zenith M1:
Total Collimation -1.046 -1.018 -0.674 -35.44 -0.09
Lookup Table      -0.801 +1.578 -1.361 -13.88 +0.89
Temperature       -0.336 -0.588   +0.420 0.00 0.00
Active Optics       +0.108 -0.303 -0.153 -3.26 -1.16
Instrument - none
Pointing Offsets - none
Group Rotation - none
Secondary Offload -0.016 -1.704 +0.420 -18.306 0.176

M2 Total     -7.010 -1.138 0.00 +114.28 +181.80

Z[ 2] =       0.00
Z[ 3] =       0.00
Z[ 4] =     -72.11
Z[ 5] =     -17.68
Z[ 6] =     -48.74
Z[ 7] =      -7.46
Z[ 8] =       2.07
Z[ 9] =       6.41
Z[10] =     -24.42
Z[11] =      25.43
=> Zernike corrections saved in file /tmp/wfscimage002610_irc.zern

Dec 11 23:41:57 lbtmu107 LBT_left_GCS: WFS: rotated Zernikes for /tmp/wfscimage002610.fits are (including 0.800000 Zscale factor)
Dec 11 23:41:57 lbtmu107 LBT_left_GCS: WFS: z(02) = 0.0000
Dec 11 23:41:57 lbtmu107 LBT_left_GCS: WFS: z(03) = 0.0000
Dec 11 23:41:57 lbtmu107 LBT_left_GCS: WFS: z(04) = -36.3098
Dec 11 23:41:57 lbtmu107 LBT_left_GCS: WFS: z(05) = -1.4674
Dec 11 23:41:57 lbtmu107 LBT_left_GCS: WFS: z(06) = -88.2413
Dec 11 23:41:57 lbtmu107 LBT_left_GCS: WFS: z(07) = 10.2085
Dec 11 23:41:57 lbtmu107 LBT_left_GCS: WFS: z(08) = -11.9956
Dec 11 23:41:57 lbtmu107 LBT_left_GCS: WFS: z(09) = 27.3076
Dec 11 23:41:57 lbtmu107 LBT_left_GCS: WFS: z(10) = -9.3190
Dec 11 23:41:57 lbtmu107 LBT_left_GCS: WFS: z(11) = 11.4936

-- JohnHill - 12 Dec 2008
Topic revision: r8 - 19 Dec 2008, JohnHill
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback